OpenStudy (kitten_is_back):

Use signed numbers to solve. A helicopter descended 1200 feet, rose 800 feet, and then descended 450 feet. What was the net gain or loss in altitude? A. lost 2450 ft B. lost 850 ft C. gained 850 ft D. gained 2450 ft

OpenStudy (welshfella):

-1200 + 800 - 450

OpenStudy (anonymous):

B.

OpenStudy (kitten_is_back):

um can you explan how its b?

OpenStudy (welshfella):

work out what i wrote

OpenStudy (anonymous):

-1200+ 800 is -400

OpenStudy (anonymous):

then add -400 + 450

OpenStudy (anonymous):

oops i meant -450

OpenStudy (kitten_is_back):



OpenStudy (welshfella):

-1200 + 800 = -400 -400-450 = -850 lost 850

OpenStudy (anonymous):

so like i said B.

OpenStudy (anonymous):

:P

OpenStudy (welshfella):

yes - but plz don't give direct answers

OpenStudy (anonymous):

..................

OpenStudy (kitten_is_back):

:d thanks for explaning it to me :D

OpenStudy (welshfella):

yw

OpenStudy (anonymous):

.................